Two electric bulbs have rating 220V – 100W. They are connected in Parallel with 110 V supply line. Find Power consumed by each bulb.
 
please explain: Will power consumed by each bulb be same?

Grade:12

1 Answers

Adarsh
6 years ago
In this type of question use power rating of bulb to calculate its resistance ..’
  220 x 220 / R = 100      (   V2/R=power )  =>  R = 484ohm …
Power by each bulb will be same as they have same potential difference ..
Total power consumption = V2/R X 2 =   50W..